A company currently pays a dividend of $1.75 per share (D0 = $1.75). It is estimated that the company's dividend will grow at a rate of 18% per year for the next 2 years, and then at a constant rate of 6% thereafter. The company's stock has a beta of 0.9, the risk-free rate is 7%, and the market risk premium is 3%. What is your estimate of the stock's current price? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

STEP 1: Calculate Cost of Equity {Ke} Ke = Rp + [Beta x Market Risk Premium) = 7+ (0.9 x 3] = 7+ 2.7 = 9.7.1. STEP2 : Calcula